OpenStudy (ohohaye):

What is the maximum height this rocket will reach? Equation: h= -5t^2+30t+10 Please show me how to do this

OpenStudy (anonymous):

max is the second coordinate of the vertex

OpenStudy (anonymous):

the first coordinate of the vertex of \[y=ax^2+bx+c\] is \(-\frac{b}{2a}\) and the second coordinate is what you get when you replace \(x\) by the first coordinate

OpenStudy (anonymous):

in your case \(a=-5,b=30\)

OpenStudy (anonymous):

did you find \[-\frac{b}{2a}=-\frac{30}{2\times (-5)}\]?

OpenStudy (ohohaye):

I'm confused

OpenStudy (anonymous):

me too what are you confused about?

OpenStudy (ohohaye):

Oh wait I get it, sorry I got lost

OpenStudy (anonymous):

ok good it is brief

OpenStudy (ohohaye):

So all I have to do is solve this and I will get the max height? \[\frac{ 30 }{ 2(-5) }\]

OpenStudy (anonymous):

you are off by a minus sign it is \[-\frac{ 30 }{ 2(-5) }\]that will get you the input yes then plug it in

OpenStudy (anonymous):

i.e. plug in \(3\) for \(t\)

OpenStudy (ohohaye):

ok, thank you